Cos3© is equivalent to
1
Expert's answer
2019-05-27T13:12:24-0400
"Cos(3x)=Cos(2x+x)=\\
=Cos(2x)Cos(x)-Sin(2x)Sin(x)=\\
=(Cos^2(x)-Sin^2(x))Cos(x)-2Sin(x)Cos(x)Sin(x)=\\
=Cos^3(x)-3Sin^2(x)Cos(x)=\\
=Cos^3(x)-3(1-Cos^2(x))Cos(x)=\\
=4Cos^3(x)-3Cos(x)"